# A Comprehensive Survey of Deep Transfer Learning for Anomaly Detection in Industrial Time Series: Methods, Applications, and Directions

### v1

@@ 4.1 Overview
Currently, deep transfer learning approaches are popular in the field of computer vision and natural language processing because of the large available datasets. This popularity is not as pronounced for industrial time series data, likely due to the lack of publicly available data and the domain-specific differences between such data, making the field a less easy target for general gains. Fortunately, in recent years a growing number of deep transfer learning approaches have been applied in industry to solve anomaly detection tasks, such as fault diagnosis (Li et al., 2022), quality management (Ma et al., 2019), manufacturing process monitoring (Tercan et al., 2018), network/software security (Rosenberg et al., 2018), and infrastructure monitoring (Pan et al., 2023). These can be mapped onto the core industrial domains of manufacturing process and infrastructure monitoring, predictive maintenance, and energy management. Table 2 presents a compact comparison of the related works using deep transfer learning approaches to solve these tasks.
@@
Figure 4 illustrates the quantity structure of the connections between industrial applications and the deep transfer learning approaches based on our literature survey. The Sankey diagram shows every path that connects the four dimensions of the methodology-problem-landscape within the surveyed literature. The broader the path is, the more papers are related to that element. The goal is to give an overview of how deep transfer learning is applied to industrial problems in the recent literature and specifically show with these four dimensions: (1) which deep transfer learning approaches are actually used in practice; (2) what the main industrial domains for time series anomaly detection are; (3) what deep transfer learning category these domains belong to; and (4) what labels are available in source and target domain.
@@
Key observations from Figure 4 are: Regarding deep transfer learning approaches, parameter transfer is much more frequently used than any other deep transfer learning approach across all surveyed industrial applications since fine-tuning a pre-trained model on target data is more straightforward to implement by taking advantage of the pre-trained model on the source dataset and usually without fundamental modification on the model architecture. It is noteworthy that instance transfer and adversarial transfer do not appear in the diagram. Apparently, these two deep transfer learning approaches are not considered the optimal choice for respective time series anomaly detection tasks in the industry. The difficulty lies in implementing and training these scarcely researched approaches in the industrial field, as indicated by the findings. Regarding industrial applications, hybrid approaches of parameter and mapping transfer can be seen in predictive maintenance. Regarding deep transfer learning categories, most industrial applications use inductive TL, indicating they focus on leveraging labeled source and target data to solve the target task, i.e., use supervised learning.
@@
CNN+LSTM
Single source
Liao et al.,2021
@@
VGG-19, ResNet
Single source
Lockner and Hopmann,2021
@@
Single source
Abdallah et al.,2021
Digital agriculture and smart manufacturing systems anomaly detection
@@
Single source
Panjapornpon et al.,2023
@@
LSTM, CNN, FCN
Single source
Liang et al.,2018
Electricity consumption anomaly detection in aluminium extrusion processes
@@
Cluster-based DAN
Single source
Simone and Amigoni,2021
@@
CNN+LSTM
Single source
Figure 4: Overview of quantity structure of transfer learning problem setting, deep transfer learning approach categories, and label availability in the surveyed industrial domains.
@@ 4.2 Manufacturing process monitoring
Manufacturing process monitoring is crucially important to ensure high-quality products and low rejection rates. For example, in injection molding machines, sensors are installed to detect molding conditions in the cavity, such as cavity pressure and temperature. These signals are used to analyze in particular the mold filling and solidification process for each produced part. Such cyclic processing data can also be seen in metal machining (cutting force signal) or joining of parts (joining force signal). Currently, parameter transfer is predominantly used for manufacturing processes (Lockner and Hopmann, 2021; Tercan et al., 2018, 2019; Lockner et al., 2022; Maschler et al., 2021; Abdallah et al., 2021, 2023; Hsieh et al., 2019; Maschler et al., 2021).
In injection molding, parameter transfer is applied to transfer the knowledge from one or more source domains to solve tasks in a target domain (Lockner and Hopmann, 2021; Tercan et al., 2019; Lockner et al., 2022). Similarly, Tercan et al. (2018) build a bridge between simulated data and real data using parameter transfer in injection molding. Maschler et al. (2021) compare different DNNs for anomaly detection tasks on metal forming datasets. Further, they propose a deep transfer learning framework aiming to transfer knowledge between tasks. However, the proposed architecture is not validated. Later, Maschler et al. (2021) apply continuous learning on the same dataset by transferring knowledge from several source tasks to a target task to train a deep learning algorithm capable of solving both source and target tasks. Abdallah et al. (2021, 2023) apply parameter transfer to monitor operation status of manufacturing testbeds with vibration sensor data. Hsieh et al. (2019) transfer knowledge across three chambers in a production line to detect anomalous time series data. Results show reduced training time and improved detection accuracy through transfer learning.
@@ 4.3 Predictive Maintenance
Predictive maintenance aims to predict the necessity of maintenance before production is negatively impacted by a failure. Tasks involve monitoring equipment to anticipate maintenance requirements (i.e., predict likely future failure) to optimize maintenance schedules (Serradilla et al., 2022). Time series anomaly detection is often used in respective systems to identify abnormal patterns or behaviors in operation that may indicate the need for maintenance, such as increasing noise, vibrations, etc.
Mao et al. (2020) use mapping transfer with a Sparse Auto-Encoder (SAE) for motor vibration anomaly detection. A transformation from the source and target space to a common latent feature space is learned by MMD to make the feature distribution of two domains as identical as possible. Similarly, Wen et al. (2019) also used mapping transfer with an SAE architecture for fault detection of rotation bearings, using an MMD regularizer to extract a common feature representation. Subsequently, they propose a new MU-Net architecture to deal with multivariate time series anomaly detection tasks (Wen and Keyes, 2019). First, they pre-train a U-Net (Ronneberger et al., 2015) on a large time series dataset for an anomaly detection task. Then, they propose a new model MU-Net, built upon U-Net, wherein each channel they can use the pre-trained U-Net through fine-tuning to transfer knowledge for multivariate time series anomaly detection.
In another application, parameter transfer is used to predict the remaining useful life for tools in manufacturing (Sun et al., 2019). An SAE network is first trained to predict the remaining useful life of a cutting tool on retrospectively acquired data in an offline process. The trained network is then transferred to production with a new tool in operation for online remaining useful life prediction. A 2D CNN-LSTM (Hochreiter and Schmidhuber, 1997) hybrid architecture for fault detection is presented. The model is trained on a fault dataset, and then parameter transfer is applied to target datasets with a different set of conditions. The result shows that transfer learning based hybrid deep learning significantly reduces the training time and is highly suitable for real-time industrial fault diagnosis in various environments. Similarly, parameter transfer is implemented to reduce the gap between different industrial environments (Zabin et al., 2023; Xu et al., 2019). Xu et al. (2019) use a stacked SAE to extract general features from source data and a digital-twin-assisted fault diagnosis approach is presented to transfer knowledge from virtual space to physical space for real-time use. Here, a DNN model is first fully trained in virtual space and then migrated to the physical space using deep transfer learning for real-time use.
The surveyed literature proves that deep transfer learning is a research field that could simplify the life cycle of predictive maintenance systems and facilitate DNN model reusability by reducing the required data and training time, helping adapt them to solve similar tasks.
@@ 4.4 Energy management
Energy management deals with systems that detect abnormal excessive consumption caused by end-users’ unusual behavior or malfunction of faulty devices or systems (Copiaco et al., 2023). The goal is to develop automatic, quick-responding, accurate, and reliable fault detection to save energy and build environmentally friendly systems. Energy anomaly detection systems monitor data during energy generation, transmission, and utilization, in order to ensure normal energy consumption. Xu et al. (2021) design a cluster-based deep adaptation layer to improve a deep adaptation network, effectively reducing the mismatch in transfer learning of spinning power consumption anomaly detection. Liang et al. (2018) successfully build an electricity consumption time series anomaly detection method in aluminum extrusion. Parameter transfer is applied to transfer domain knowledge from another data-sufficient domain. They also find it unnecessary when the target data is sufficient because transferring knowledge decreases prediction accuracy.
@@ 4.5 Infrastructure facilities monitoring
Infrastructure facilities monitoring refers to monitoring and maintaining the conditions of infrastructure facilities, such as bridges, buildings, and networks. This can include detecting potential issues or failures. The goal is to minimize the impact of failures on the public or the environment. Pan et al. (2023) apply parameter transfer to make full use of the similarity of the anomalous patterns across different bridges and transfer the knowledge obtained by a CNN model to a small part of target data, achieving high accuracy anomaly detection across bridges. Dhillon and Haque (2020) present a parameter transfer approach towards building a network intrusion detection system based on CNN and LSTM.
@@ 5.1 Challenges
Label availability
Deep transfer learning is built upon deep learning, which usually requires a large amount of labeled data, the more data a model has available for learning, the better it can generalize to new examples. In real-world industrial time series anomaly detection tasks, collecting data is probably easy, but collecting labels is much more expensive and time-consuming, sometimes prohibitively so, leading to the unavailability of sufficient labeled data. Self-supervised learning can be used to re-label a large amount of unlabeled data and thus facilitate transfer learning process. Thus, anomaly detection models usually need to learn in an unsupervised or semi-supervised mode (Goldstein and Uchida, 2016).
Deep learning for imbalanced data
Even if the labels can be collected, anomalies can be extremely rare by design, which poses the risk of training with extremely imbalanced data. A practical problem for anomaly detection in industry is the extremely imbalanced data distribution, in which normal samples dominate in data and abnormal samples only share a small percentage in the whole dataset. Prior research has proven that the effect of class imbalance on classification performance by using deep learning is detrimental (Buda et al., 2018). However, most research studies still ignore such problems, which can result in poor performance regarding the minority class, i.e., abnormal data are misclassified as normal.
Missing relevant data
Another problem is missing relevant data, i.e, some information that has a significant effect on the process from case to case is not even recorded or is too complex to record (i.e. part geometry, machine geometry, or environmental conditions in injection molding processes).
@@
Domain shifts lie at the heart of the deep transfer learning problem, but the dynamic changes in many industrial processes, up to an apparent dissimilarity of source and target data, make the transfer learning task particularly challenging.
@@
The general effectiveness of deep transfer learning is limited by the difficulty of determining which knowledge or to what extent the knowledge should be transferred from source to target task. Unlike natural language processing, pretraining a language model on a large corpus of text data can help the model learn the statistical patterns and semantic and syntactic representations of words and sentences, which can be used for new natural language processing tasks with a few data. For industrial time series, due to data privacy, large available public datasets usually do not exist, or they cannot be used even because of a large domain gap between different datasets and tasks. In this case, transferring all of the knowledge may not be beneficial, as it may be irrelevant. In the worst case, this can lead to negative transfer (Pan and Yang, 2010; Smith-Jentsch et al., 2001), in which the extracted knowledge harms the new task-learning. This requires assessing how source and target tasks are related, carefully selecting the knowledge to be transferred, and selecting the proper means to implement this transfer.

### v2

@@ 4.1 Overview
Deep transfer learning techniques have gained prominence in computer vision and natural language processing, primarily due to the abundance of available datasets. However, their adoption in the context of industrial time series data has been comparatively limited. This hesitancy can be attributed to the limited public availability of such datasets and the unique domain-specific characteristics they possess, which complicate generalized advancements. Encouragingly, there has been a recent uptick in the application of deep transfer learning for anomaly detection within the industry such as fault diagnosis [105], quality management [106], manufacturing process monitoring [85], network/software security [107], and infrastructure monitoring [108]. These can be mapped onto the core industrial domains of manufacturing process and infrastructure monitoring, predictive maintenance, and energy management. Table 2 presents a compact comparison of the related works using deep transfer learning approaches to solve these tasks.
@@
Fig. 7 illustrates the Sankey diagram of the connections between industrial applications and the deep transfer learning approaches based on our literature survey. The diagram shows every path that connects the four dimensions of the methodology-problem-landscape within the surveyed literature. The broader the path is, the more papers are related to the linked topics. The goal is to give an overview of how deep transfer learning is applied to industrial problems in the recent literature and specifically show with these four dimensions: (1) which deep transfer learning approaches are actually used in practice; (2) what the main industrial domains for time series anomaly detection are; (3) what deep transfer learning category these domains belong to; (4) what labels are available in source and target domain.
@@
Key observations from Fig. 7 are: (1) parameter transfer is much more frequently used than any other deep transfer learning approach across all surveyed industrial applications since fine-tuning a pre-trained model on target data is more straightforward to implement by taking advantage of the pre-trained model on the source dataset and usually without fundamental modification on the model architecture. It is noteworthy that instance transfer and adversarial transfer do not appear in the diagram. Apparently, these two deep transfer learning approaches are not considered effective in time series anomaly detection tasks in industry. The difficulty lies in implementing and training these scarcely researched approaches in the industrial field, as indicated by the findings. (2) Hybrid approaches of parameter and mapping transfer can be seen in predictive maintenance. (3) Most industrial applications use inductive transfer learning, indicating they focus on leveraging labeled source and target data to solve the target task, i.e., use supervised learning.
Figure 7: Overview of Sankey diagram of transfer learning problem setting, deep transfer learning approach categories, and label availability in the surveyed industrial domains.
@@
CNN, LSTM
Single
[9]
@@
VGG, ResNet
Single
[10]
@@
Single
[20]
Agriculture/manufacturing systems anomaly detection
@@
Single
[118]
Industrial control systems anomaly detection
Manufacturing process monitoring
Parameter transfer
Inductive
Auto-encoder
Single
[21]
@@
LSTM, CNN
Single
[119]
Electricity consumption anomaly detection
@@
DAN
Single
[122]
@@
CNN, LSTM
Single
[17]
Building occupation detection
Infrastructure facilities monitoring
Parameter transfer
Inductive
CNN
Single
[18]
Building occupation detection
Infrastructure facilities monitoring
Parameter transfer
Inductive
CNN
Single
@@ 4.2 Manufacturing process monitoring
Manufacturing process monitoring is crucial to ensure high-quality products and low rejection rates. For example, in injection molding machines, sensors are installed to detect molding conditions in the cavities, such as cavity pressure and temperature. These signals are used to analyze particularly the mold filling and solidification process for each produced part. Such cyclic processing data can also be seen in metal machining (cutting force signal) or joining of parts (joining force signal). Currently, parameter transfer is predominantly used for manufacturing processes [10, 85, 113, 114, 11, 20, 116, 123, 117].
Park et al. propose a transfer learning technique to detect time series anomalies for different industrial control systems [118]. First, they apply principal components analysis to reduce the dimension of source and target data. A DNN model is then trained on the compressed source data, and after a reasonable mapping algorithm is adopted to map the features of source to target domain, the pre-trained model is further trained on the target data. The model achieves good performance even when a model is retrained with only a proportion of target data. For the experiments, they only test on two comparatively larger datasets and fail to show that the transferred model performs better than the one without retraining for one dataset. Further investigation needs to explain the negative transfer. Additionally, even when they only take a small proportion of target data for transfer learning purposes, the sample size still exceeds $5000$, which exceeds most industrial applications. Abdallah et al. apply parameter transfer to monitor the operation status of manufacturing testbeds with vibration sensor data [20, 116]. Hsieh et al. transfer knowledge across three chambers in a production line to detect anomalous time series data [123]. Results show reduced training time and improved detection accuracy through transfer learning. In injection molding, parameter transfer is applied to transfer the knowledge from one or more source domains to solve tasks in a target domain [10, 113, 114]. Specifically, they employ simple fully connected neural networks and transfer knowledge from one product to another by freezing the first few layers and fine-tuning only the last few layers. Instead of evaluating the time series data directly from sensors, they represent the industrial process by the parameters of the machine settings. However, they can still provide useful insight for the case of the time-series data. Tercan et al. build a bridge between simulated data and real data using parameter transfer in injection molding [85]. Here, a fully connected neural network is trained on simulated data and then partially or fully reused to further train on real data. Results show that the transferred model performs better than a network trained from scratch on real experimental data. In manufacturing processes, a simulation model/process hence can play a critical role, but deeper analysis is needed to further understand and reduce the gap between simulation and real data. Additionally, Lockner et al. explore the impact of transfer learning with varying amounts of source data and assess how performance is influenced by different configurations of frozen layers [114]. Maschler et al. compare different DNNs for anomaly detection tasks on metal forming datasets [11]. Further, they propose a deep transfer learning framework aiming to transfer knowledge between tasks. However, the proposed architecture is not validated by experiments. Later, Maschler et al. apply continuous learning on the same dataset by transferring knowledge from several source tasks to a target task to train a deep learning algorithm capable of solving both source and target tasks [117]. Specifically, they use regularization approaches using altered loss functions to solve related tasks that appear best suited.
@@ 4.3 Predictive maintenance
Predictive maintenance aims to predict the necessity of maintenance before production is negatively impacted by a failure. Tasks involve monitoring equipment to anticipate maintenance requirements (i.e., predict probable future failure) to optimize maintenance schedules [124]. Time series anomaly detection is often used in respective systems to identify abnormal behaviors in operation that may indicate the need for maintenance, such as increasing noise, vibrations, etc.
Mao et al. use mapping transfer with a Sparse Auto-Encoder (SAE) for motor vibration anomaly detection [14]. A transformation from the source and target data to a common latent feature space is learned with MMD loss to make the feature distribution of two domains as identical as possible. Similarly, Wen et al. also use mapping transfer with an SAE architecture for fault detection of rotation bearings, using an MMD regularizer to extract a common feature representation [125]. Subsequently, they propose a new MU-Net architecture to detect multivariate time series anomalies [12]. First, they pre-train a U-Net [126] on a large time series dataset for an anomaly detection task. Then, they propose a new model MU-Net, which is built upon U-Net. In MU-Net, each channel can leverage a pre-trained U-Net through fine-tuning to transfer knowledge for multivariate time series anomaly detection.
In a different application, parameter transfer is used to predict the remaining useful life for tools in manufacturing [127]. An SAE network is first trained to predict the remaining useful life of a cutting tool on retrospectively acquired data in an offline process. The trained network is then transferred to production with a new tool for online remaining useful life prediction. The result shows that transfer-learning based hybrid deep learning significantly reduces the training time and is highly suitable for real-time industrial fault diagnosis/prediction in various environments. Similarly, parameter transfer is implemented to reduce the gap between different industrial environments [112, 13]. Xu et al. use a stacked SAE to extract general features from source data and a digital-twin-assisted fault diagnosis approach is presented to transfer knowledge from virtual space to physical space for real-time use [13]. Here, a DNN model is first fully trained in virtual space and then migrated to the physical space using parameter transfer for real-time use.
The above-mentioned literature proves that deep transfer learning is a research field that could simplify the life cycle of predictive maintenance systems and facilitate DNN model reusability by reducing the required data and training time, helping adapt them to solve similar tasks.
@@ 4.4 Energy management
Energy management deals with systems that detect abnormal excessive consumption caused by end-users’ unusual behavior or malfunction of faulty devices or systems [120]. The goal is to develop automatic, quick-responding, accurate, and reliable fault detection to save energy and build environmentally friendly systems. Energy anomaly detection systems monitor data during energy generation, transmission, and utilization, to ensure normal energy consumption.
Xu et al. design a cluster-based deep adaptation layer to improve a deep adaptation network, effectively reducing the mismatch in transfer learning of spinning power consumption anomaly detection [121]. The basic architecture consists of five convolutional layers and three fully connected layers. The weight parameters of the convolutional layers are shared between source and target domains. The cluster-based deep adaptation layer is designed across the feature layers of two networks to cluster feature representations of the source and target domains respectively. The proposed method shows superiority over fine-tuning and DAN because the adaptation layer can minimize the distance between the nearest neighbor clusters across the source and target domains to match the most similar distribution of feature representations. It is important to note that the anomalies are defined and tagged by human experts as different types, thus the problem becomes a classification task. However, in real-world industrial applications, it’s almost impossible to enumerate unknown anomaly types because of the highly dynamic environment. Liang et al. successfully build an electricity consumption time series anomaly detection method in aluminum extrusion [119]. Parameter transfer is applied to transfer domain knowledge from another data-sufficient domain. First, they train on sufficient extruding machine data in an unsupervised way and then use only a few data samples from different extruding machines to adapt the model by transfer learning. It is important to note that when the target data is already sufficient, transferring knowledge can be detrimental as it can decrease prediction accuracy on the final task. Copiaco et al. aims to detect anomalies for building energy consumption via transfer learning from pre-trained CNN models [128]. First, they convert 1D time series signals to 2D image representations. These serve as inputs for pre-trained vision models to capture inherent spatially invariant features. In the end, a SVM is applied to classify anomaly types. The SVM classifier obtained optimal results when operating upon a pre-trained ALexNet model with normalized grayscale graphical representations. However, a deeper discussion regarding the effect of the different pre-trained models is not presented. Additionally, converting 1D time series to 2D images by creating a matrix representation of the sensor readings may lead to information loss during the transformation process, which should be further investigated.
@@ 4.5 Infrastructure facilities monitoring
Infrastructure facilities monitoring refers to monitoring and maintaining the conditions of infrastructure facilities, such as bridges, buildings[129], and networks. This can include detecting potential problems or failures. The goal is to minimize the impact of failures on the public or the environment. This application commonly uses parameter transfer to transfer knowledge from facility to facility to take advantage of similar data and tasks.
Dhillon et al. present a parameter transfer approach towards building a network intrusion detection system based on CNN and LSTM [22]. Specifically, They extract and learn patterns by mapping the input data into a lower dimensional representation by convolutional layers. Then, they employ the LSTM layer to enhance learning and recognizing patterns across time. In the end, a fully connected layer is used as a classifier to predict normal and malicious data. To do the parameter transfer, they reuse the model architecture and freeze most weight parameters for the target domain so that they do not need a large training dataset to retrain the model. However, they do not mention implementation details, like which layers are frozen in the transfer learning stage. Observing how transfer learning performs with different frozen layers would be interesting. Pan et al. apply parameter transfer to fully use the similarity of the anomalous patterns across different bridges [108]. They train a CNN model on one bridge data, then transfer the knowledge obtained by the CNN model to a small part of the target data. They update the last three fully connected layers while keeping the convolutional layers intact. The experimental results show transfer learning achieves higher accuracy anomaly detection across bridges. Weber et al. takes advantage of simulation data by training on synthetic environmental data, then fine-tunes the pre-trained model and transfers the knowledge from simulation data for real-time online building occupancy detection [17]. Although the results show the effectiveness of transfer learning, the availability, and reliability of the simulation data for other industrial applications is still an open issue. Sayed et al. adapt parameter transfer using pre-trained CNN models, such as AlexNet and GoogLeNet, pre-trained on ImageNet [18]. The pre-trained model is then further used for downstream tasks. The results show the pre-trained models outperform their customed CNN model, which is not pre-trained. However, it is important to note that the transfer effect may not be entirely convincing due to the fact that the customer CNN model is not pre-trained on the same dataset.
4.6 Application-independent considerations
[…]
Data scarcity, as well as domain shift, stand out as the two main common problems independent of the industrial field of application. The same problems have originally prompted the use of transfer learning in general, and respectively, general techniques are applied widely across domains. Regarding data scarcity, this un-surprisingly involves leveraging pre-trained models as a starting point for further training. Regarding domain shift, mapping transfer and parameter transfer are the most often-used approaches. Unlike parameter transfer, mapping transfer incorporates the source and target data in the training process. Instance transfer and domain-adversarial transfer learning were not employed in the surveyed literature – researchers seem to not see huge value in these methods for the surveyed fields.
Another common aspect across time series anomaly detection applications is the choices of model architecture to facilitate the training process by capturing temporal dependencies and recognizing patterns over varying time scales: Favourite architectures include CNNs, LSTMs, and auto-encoders that have sets of assumptions (inductive biases) about the data they analyze that make them excel in understanding the sequential nature of time series data. CNNs assume local (in time) connectivity, stationary statistics, and hierarchical structure, and induce certain translation-invariance. LSTMs are still given preference in many applications over the more modern deep learning architecture of choice for sequence learning, the transformer. The reason is their stronger inductive bias, leading to less data (and compute) hunger. Both CNN and LSTM networks can be built as e.g. classifiers, but also auto-encoders. These latter architectures have the advantage of learning low-dimensional representations of the high-dimensional time series with minimal loss of signal in an unsupervised way. The analysis of the data in the low-dimensional latent space facilitates anomaly detection. The concrete choice of architecture does not depend on the field of application but on the data and task, and thus the most suitable inductive bias.
As an interim conclusion, the most striking application-independent finding is that across the surveyed literature, predominantly simple, tried-and-tested design patterns for transfer learning are used in industry. The field of deep transfer learning would offer a much wider variety of approaches.
@@ 5 Discussion
5.1 Potential
The automation of industrial process monitoring stands as a transformative step toward increasing efficiency and optimizing quality. While standard deep learning training is sufficient in discerning intricate patterns from vast datasets, its application in the dynamic industrial landscape is not without challenges. Chief among them is the impracticality of continuously obtaining large-scale labeled data to train models afresh for every nuanced variation in processes. Deep transfer learning has shown promise with its adaptive capabilities. By mitigating the need for extensive labeled data and eliminating the necessity to train models from scratch for every distinct setup. However, adopting deep transfer learning beyond simple parameter transfer is still a challenge.
@@ 5.2 Challenges
Different from the i.i.d assumption in most machine learning problem settings, many industrial processes suffer from substantial domain shift due to dynamic changes in industrial settings, e.g., change of products or measuring sensors. Domain shift lies at the heart of the deep transfer learning problem. Particularly, the dynamic changes in many industrial processes, up to an apparent dissimilarity of source and target data, make the transfer learning task particularly challenging. Here we list the key challenges associated with domain shift:
  • Covariate shift occurs when the marginal distribution of the features changes from source domain to target domain. The distribution mismatch poses challenges in transferring the knowledge from source to target domain.
  • Concept shift refers to the changes in the relationships between features and labels. The relationship can change from source domain to target domain, leading to bias and error in the model.
  • Label shift refers to the label distribution in the target domain that can be different from the source domain, whether the marginal distribution changes or not.
Label availability and reliability
Deep transfer learning is built upon deep learning, which usually requires a large amount of labeled data, the more data a model has available for training, the better it can generalize to new examples. In real-world industrial time series anomaly detection tasks, collecting data is probably easy, but collecting labels is much more expensive and time-consuming, sometimes prohibitively so, leading to the unavailability of sufficient labeled data. Self-supervised learning can be used to re-label a large amount of unlabeled data and thus anomaly detection models usually need to learn in an unsupervised or semi-supervised mode [130]. In industrial cases, another significant concern is to ensure the data quality. Due to the high cost associated with obtaining reliable and precise labels, usually self-supervised learning is applied to create pseudo labels or relabel the unlabeled data, thus facilitating the transfer learning process. Additionally, a data-centric process with humans in the loop can be involved in improving label reliability. However, unreliable labels can still affect the transfer learning training process.
Missing relevant data information
Missing relevant data poses significant challenges for transfer learning since it can affect the model’s ability to generalize and transfer knowledge from source to target domain.
  • Imbalanced data: Even if the labels can be collected, anomalies can be extremely rare by design, which poses the risk of training with extremely imbalanced data. A practical problem for anomaly detection in industry is the extremely imbalanced data distribution, in which normal samples dominate in data and abnormal samples only share a small percentage in the whole dataset. Prior research has proven that the effect of class imbalance on classification performance by using deep learning is detrimental [131]. However, most research studies still ignore such problems, which can result in poor performance regarding the minority class, i.e., abnormal data are misclassified as normal.
  • Information loss: missing data can lead to lost important features. For example, some information that has a significant effect on the process from case to case is not even recorded or is too complex to record (i.e. part geometry, machine geometry, or environmental conditions in injection molding processes).
Various approaches have been developed to address these challenges to reduce the domain gap between the source and target domains, aiming at mitigating domain shift. These techniques involve domain generalization, contrastive learning, and adversarial examples. The domain shift problem is far from being solved. To tackle this problem, transfer learning requires a deep understanding of the target data’s characteristics and appropriate transferable strategies to effectively bridge the gap between the source and target domains.
@@
The general effectiveness of deep transfer learning is limited by the difficulty of determining which knowledge or to what extent the knowledge should be transferred from source to target task. Unlike natural language processing, pre-training a language model on a large corpus of text data can help the model learn the statistical patterns and semantic and syntactic representations of words and sentences, which can be used for new natural language processing tasks with a few or even without data. Due to data privacy, large available public datasets usually do not exist for industrial time series, or they can not be used because of a large domain gap between different datasets and tasks. In this case, transferring all of the knowledge may not be beneficial, as it may be irrelevant. In the worst case, this can lead to negative transfer [28, 132], in which the extracted knowledge harms the new task-learning. This requires assessing how source and target tasks are related, carefully selecting the knowledge to be transferred, and selecting the proper means to implement this transfer. Glorot et al. attempt to analyze and quantify the gained knowledge from source to target domain [133]. For example, they define transfer error, transfer loss, transfer ratio, and in-domain ratio, which provide metrics to interpret the transferring performance.